A hypothetical demand and supply functions are as follows: 300 = Qd + 1.5P and P = 25 + Qs.
a. Calculate the equilibrium price and equilibrium quantity.
b. What would be the new quantity demanded and supplied if the price reduce to 65.00?
1
Expert's answer
2019-12-12T06:06:30-0500

1. Equilibrium price and quantity:

300=Q+1.5×(25+Q)

300=2.5Q+37.5

2.5Q=262.5

Q=105

P=130

2. If price is 65

Qd=300-1.5×65=202.5

Qs=65-25=40